Quantitatively and qualitatively discuss what happens to the spring constant when springs are placed in a series or parallel configuration.
Answer #1


when the springs are connected in series

net spring constant is K

1/K = (1/K1)+(1/K2)

K = (K1*K2)/(K1+K2)

in series connection ,net spring constant is less than any other individual spring constants

IN parallel combination

K = K1+K2


In parallel combination the net spring constant is greater than any other individual spring constants

2

2